# Adaptive Feature Selection for No-Reference Image Quality Assessment by Mitigating Semantic Noise Sensitivity

### v2

@@ Abstract
The current state-of-the-art No-Reference Image Quality Assessment (NR-IQA) methods typically rely on feature extraction from upstream semantic backbone networks, assuming that all extracted features are relevant. However, we make a key observation that not all features are beneficial, and some may even be harmful, necessitating careful selection. Empirically, we find that many image pairs with small feature spatial distances can have vastly different quality scores, indicating that the extracted features may contain a significant amount of quality-irrelevant noise. To address this issue, we propose a Quality-Aware Feature Matching IQA Metric (QFM-IQM) that employs an adversarial perspective to remove harmful semantic noise features from the upstream task. Specifically, QFM-IQM enhances the semantic noise distinguish capabilities by matching image pairs with similar quality scores but varying semantic features as adversarial semantic noise and adaptively adjusting the upstream task’s features by reducing sensitivity to adversarial noise perturbation. Furthermore, we utilize a distillation framework to expand the dataset and improve the model’s generalization ability. Our approach achieves superior performance to the state-of-the-art NR-IQA methods on eight standard IQA datasets.
@@ 1 Introduction
However, we empirically find that many image pairs with small feature spatial distances have vastly different quality scores (see 4.6). We attribute this result to feature aliasing, where semantic noise unrelated to quality confuses the quality-aware features. Specifically, due to the sensitivity of upstream backbone networks to semantic features (Zhao et al., 2023; Zhang et al., 2023), they tend to extract similar features for semantically similar image pairs, despite these pairs having distinctly different quality scores. This may mislead the network to overly focus on semantic details rather than true indicators of image quality, resulting in poor quality prediction. We provide an example in Fig.1(b) to illustrate this issue. For a pair of images with obvious quality differences, the baseline overly focuses on similar semantic information in the distorted image (e.g., the person in the yellow box), while ignoring some distorted areas in the red box (e.g., the racket’s movement and the flower’s overexposure), causing the model to be misled into making a close quality predictions. These findings motivate us to explore a critical problem: how to distinguish images with similar semantic information but different quality scores.
We believe the key to addressing this challenge is to reduce the model’s sensitivity to semantic noise. Inspired by adversarial learning (Goodfellow et al., 2014; Madry et al., 2017), which aims to learn adversarial noise that maximizes prediction loss, and reduces the model’s sensitivity to this noise by minimizing the model’s loss when facing noisy inputs. In this paper, we treat samples with similar quality scores but significantly different semantic features as adversarial noise to maximally mislead the model’s quality prediction. This approach is based on our hypothesis that the model is highly sensitive to semantics. Therefore, introducing maximal semantic perturbation can greatly impact its quality predictions, while similar quality scores ensure that the misprediction is mainly influenced by the semantic perturbation rather than the quality perturbation.
@@
Figure 2: The overview of our QFM-IQM.We begin by extracting image semantic-related features from the student network. Simultaneously, we indirectly augment our training set by introducing a distillation framework and storing the image features and labels in pairs in temporary memory (Sec. 3.4). We use Semantic Noisy Matching to select features in the temporary memory with similar quality scores but different semantic features as noise disturbance (Sec. 3.5) and add them to our original features as decoder input to predict image quality, forcing our model to learn to focus on quality-aware features (Sec. 3.6). During inference, scores are directly predicted using the encoder and decoder, while the above three modules are discarded to avoid any additional computational overhead.
@@ 2 Related Works
Image Quality Assessment (IQA) (Hu et al., 2021) is broadly classified into full-reference (FR) (Cao et al., 2022), reduced-reference (RR) (Tao et al., 2009), and no-reference (NR) methods (Zhang et al., 2021). NR-IQA assesses quality independently, offering significant application potential.
@@ 2.1 NR-IQA with Vision Transformer
Vision Transformer (ViT) (Dosovitskiy et al., 2021) showed promising results on several downstream vision tasks. There were mainly two types of ViT-based NR-IQA methods, including hybrid Transformer  (Golestaneh et al., 2022) and pure ViT-based Transformer  (Ke et al., 2021b; Yang et al., 2022; You & Korhonen, 2021). The hybrid architecture generally combined the CNNs with the Transformer, which was responsible for the local and long-range feature characterization, respectively. For instance, (Golestaneh et al., 2022) proposed to use the multi-scale features extracted from ResNet-50, which were fed to the transformer encoder to produce a non-local representation of the image. (Ke et al., 2021b) designed a multi-scale ViT-based IQA model to handle the arbitrary size of input images. Such a feature was initially designed to describe the image content, and thus the preserved features were mainly related to the higher-level visual abstractions which were not adequate in characterizing the quality-aware
@@ 2.2 NR-IQA with Contrastive Learning
A new learning paradigm, contrastive learning, emerged for learning discriminative representations among samples (Caron et al., 2020; He et al., 2020). This method aimed to learn an embedding space where similar samples were attracted, and dissimilar ones were repelled (Jaiswal et al., 2020; Zhao et al., 2023). Recently, contrastive learning in IQA primarily focused on unsupervised approaches, as seen in key studies like QPT (Zhao et al., 2023), which introduced a self-supervised method with quality-aware contrastive learning for BIQA. Re-IQA (Saha et al., 2023) employed a mixture of experts to train separate encoders for content and quality features, and CONTRIQUE (Madhusudana et al., 2022) used a deep CNN with a contrastive pairwise objective for IQA. These methods achieved promising results by leveraging self-supervised training, but the cost of pre-training models was often high. In contrast to contrastive learning aimed at constructing discriminative embedding spaces, we leveraged adversarial learning to reduce the model’s sensitivity to semantic noise. This enhanced the model’s ability to distinguish between images that were semantically similar but had different quality scores.
@@ 3.1 Preliminaries
In the context of No-Reference Image Quality Assessment (NR-IQA), we define some common notations. We use bold formatting to denote vectors (e.g., $\boldsymbol{x}$, $\boldsymbol{y}$) and matrices (e.g., $\boldsymbol{X}$, $\boldsymbol{Y}$). Specifically, $\boldsymbol{F}$ denotes the feature map of the network output, $\mathcal{F}$ denotes the vision transformer’s class token (CLS token), and Y denotes the quality score.
@@ 3.2 Overview
In this paper, we present Quality-Aware Feature Matching Image Quality Metric (QFM-IQM), a novel framework that can effectively distinguish between images with similar semantic features but different quality scores. As depicted in Fig. 2, QFM-IQM seamlessly integrates three main components: Semantic Noise Matching (SNM), Quality Consistency Constraint (QCC) and Distillation Label Extension (DLE). Initially, the input distorted image is fed into a transformer encoder to obtain feature $\boldsymbol{F}_{o}\in\mathbb{R}^{(M+1)\times D}$ from the final layer. Subsequently, the SNM module matches these features with those that have similar quality scores but different semantics (Sec. 3.5). Next, the matched semantic noise is introduced as feature perturbation into feature $\boldsymbol{F}_{o}$ to obtain $\hat{\boldsymbol{F}_{o}}$ and the QCC module constraints between the quality prediction of the $\boldsymbol{F}_{o}$ and $\hat{\boldsymbol{F}_{o}}$ to remain consistent, forcing the model to be robust to quality-unrelated semantic noise during adversarial learning (Sec. 3.6). Additionally, the DLE module uses knowledge distillation to generate pseudo-labels for unlabeled samples, thereby enriching the adversarial learning dataset of QCC (Sec. 3.4). These components collectively contribute to a more effective and accurate NR-IQA process. It’s worth noting that during the inference process, scores are predicted directly using the encoder and decoder, discarding the above three modules, and avoiding any additional computational overhead.
@@ 3.3 Architecture Design
Transformer Decoder. Previous works (Qin et al., 2023) found that CLS token cannot build an optimal representation for image quality. we utilize a quality-aware decoder that further interprets the CLS token. The decoder uses Multi-Head Cross-Attention (MHCA) with a set of queries $Q_{d}$, interacting with encoder outputs to calculate the final quality score $\boldsymbol{\hat{Y}}$. This method improves the learning and generalization capabilities of our NR-IQA model.
@@ 3.5 Semantic Noise Matching
Semantic Noise Feature Matching. This process involves selecting images with similar quality but different semantics to use as semantic noise features. Then, these noise features are employed as feature perturbations to disrupt semantics while largely preserving quality information. Finally, under the constraint of quality consistency (referenced in Section 3.6), this effectively reduces sensitivity to semantic noise. Fig. 3 overviews our Semantic Noise Matching (SNM) module. For a mini-batch of distorted input images, Temporary Memory $\mathcal{A}$ and $\mathcal{B}$ store collections of features and labels, then QFM-IQM leverages Semantic Noise Matching to select noise samples that have similar quality scores but different semantic features from both labeled and unlabeled datasets, as depicted in Fig. 3. Specifically, the process begins by flattening the feature $\boldsymbol{F}_{o}$ of the input image, obtained from the encoder, into a feature vector $\boldsymbol{f}_{o}$. Subsequently, another feature vector $\hat{\boldsymbol{f}}_{o}$ is matched from the Temporary Memory $\mathcal{A}$ or $\mathcal{B}$ based on the matching strategy. The evaluation metric for the matching strategy is formulated as the product of feature similarity and score distance. Feature similarity is calculated using cosine distance, while score distance is determined using the $\mathcal{L}_{2}$ norm. The mathematical expression is as follows:
$\mathcal{S}((\boldsymbol{{f}}_{o},\rm{{y}}_{o}),(\boldsymbol{\hat{{f}}_{o}},\rm{\hat{{y}}_{o}}))=\frac{\boldsymbol{{f}}_{o}\cdot\hat{\boldsymbol{f}}_{o}}{\|\boldsymbol{{f}}_{o}\|\cdot\|\boldsymbol{\hat{{f}}_{o}}\|}\cdot\|\rm{{y}}_{o}-\rm{\hat{{y}}_{o}}\|.$ (1)
@@ 3.6 Quality Consistency Constraint
Existing studies have not addressed the challenge of differentiating between images with similar semantic features but different quality scores and eliminating quality-irrelevant features from high-level pre-trained features. To address this, we introduce the Quality Consistency Constraint (QCC) module to isolate quality-related attributes from an image’s semantic content. In contrast to previous contrastive learning approaches, this paper introduces a straightforward yet effective adversarial method to enhance a model’s robustness to features unrelated to image quality. This is accomplished by strengthening the consistency of the model’s predictions for image quality before and after introducing noise to distorted images. This method reduces the model’s sensitivity to semantic noise and compels it to exhibit robustness to subtle variations, ultimately increasing the separation between feature representations of noise samples. Concretely, we initially extract the CLS tokens of the minimum K features matched in the SNM module and incorporate them as feature perturbations into the original CLS token. Since we perform feature selection twice in labeled dataset $\mathcal{A}$ and unlabelled dataset $\mathcal{B}$, the expression of this process is:
[…]
@@ 4.4 Generalization Capability Validation
To assess the generalization capabilities of QFM-IQM, we conduct a series of cross-dataset validation experiments. In these tests, QFM-IQM is trained on one dataset and then evaluated on different datasets without any adjustments to its parameters. The results, presented in Table 2, demonstrate QFM-IQM’s superior performance over SOTA models across six datasets. This includes notable gains on the LIVEC dataset and strong results on the KonIQ dataset. These results, benefiting from the noise-resistant training provided by the SNM and QCC, as well as the extensive semantic noise of various distortions provided by the DLE, underscore QFM-IQM’s exceptional ability to generalize.
@@ 4.5 Ablation Study
The core idea of our QFM-IQM is to improve the model’s ability to distinguish subtle quality differences by enhancing its robustness to semantic noise. Therefore, we explore the impact of Quality Consistency Constraint (QCC) operation on the model when compared with different datasets in Table 3. Let $\mathcal{A}^{QCC}$ denotes the QCC operation only compare with the labeled dataset $\mathcal{A}$, and $\mathcal{B}^{QCC}$ denotes the QCC operation only compare with the unlabeled dataset $\mathcal{B}$. In the ablation experiments on QCC, we use Semantic Noise Matching (SNM) by default.
@@
Figure 5: Activation maps of DEIQT  (Qin et al., 2023) and QFM-IQM using Grad-CAM  (Selvaraju et al., 2017) on authentic dataset LIVEC. The scores in the figure represent Mean Opinion Scores (MOS). Our model pays more attention to image distortion regions (red boxes) and is less sensitive to semantic noise (yellow boxes), and correspondingly our image quality prediction ability is closer to the true value. Rows 1-3 represent input images, CAMs from DEIQT  (Qin et al., 2023), and CAMs from QFM-IQM, respectively.
Effective on SNM and QCC. We broaden our investigation to include ablation studies on SNM and QCC mechanisms. Specifically, we undertake two sets of SNM ablation experiments: the first set, ${\text{SNM}}^{F}$, matches pairs of samples dissimilar in features, while the second set, ${\text{SNM}}^{Q}$, matches pairs similar in quality. Additionally, an extra ablation study without using the QCC involves matching noise samples $(x_{1},y_{1})$ and $(x_{2},y_{2})$ to a given pair $(x,y)$, where $y$ is the quality label, adjusting the quality labels by a specific ratio ($\beta_{1}=\beta_{2}=0.1$) to incorporate noise sample labels, formulated as $y=(1-\beta_{1}-\beta_{2})y+{\beta_{1}}{y_{1}}+{\beta_{2}}{y_{2}}$.
@@
Table 5: Ablation experiments on the DLE module using unlabeled dataset selection on the LIVEC and TID2013 datasets.

### v3

@@ Abstract
The current state-of-the-art No-Reference Image Quality Assessment (NR-IQA) methods typically rely on feature extraction from upstream semantic backbone networks, assuming that all extracted features are relevant. However, we make a key observation that not all features are beneficial, and some may even be harmful, necessitating careful selection. Empirically, we find that many image pairs with small feature spatial distances can have vastly different quality scores, indicating that the extracted features may contain quality-irrelevant noise. To address this issue, we propose a Quality-Aware Feature Matching IQA Metric (QFM-IQM) that employs an adversarial perspective to remove harmful semantic noise features from the upstream task. Specifically, QFM-IQM enhances the semantic noise distinguish capabilities by matching image pairs with similar quality scores but varying semantic features as adversarial semantic noise and adaptively adjusting the upstream task’s features by reducing sensitivity to adversarial noise perturbation. Furthermore, we utilize a distillation framework to expand the dataset and improve the model’s generalization ability. Extensive experiments conducted on eight standard IQA datasets have demonstrated the effectiveness of our proposed QFM-IQM.
@@ 1 Introduction
However, we empirically find that many image pairs with small feature spatial distances have vastly different quality scores (Sec. 4.6). We attribute this result to feature aliasing, where semantic noise unrelated to quality confuses the quality-aware features. Specifically, due to the sensitivity of upstream backbone networks to semantic features (Zhao et al., 2023; Zhang et al., 2023), they tend to extract similar features for semantically similar image pairs, despite these pairs having distinctly different quality scores. This may mislead the network to overly focus on semantic details rather than true indicators of image quality, resulting in poor quality prediction. We provide an example in Fig.1(b) to illustrate this issue. For a pair of images with obvious quality differences, the baseline overly focuses on similar semantic information in the distorted image (e.g., the person in the yellow box), while ignoring some distorted areas in the red box (e.g., the racket’s movement and the flower’s overexposure), causing the model to be misled into making a close quality predictions. These findings motivate us to explore a critical problem: how to distinguish images with similar semantic information but different quality scores.
We believe the key to addressing this challenge is to reduce the model’s sensitivity to semantic noise. Inspired by adversarial learning (Goodfellow et al., 2014; Madry et al., 2017), which aims to learn adversarial noise that maximizes prediction loss, and reduces the model’s sensitivity to this noise by minimizing the model’s loss when facing noisy inputs. In this paper, we treat samples with similar quality scores but significantly different semantic features compared to the input image as adversarial noise. This approach is based on our hypothesis that the pre-trained upstream backbone is highly sensitive to semantics. Therefore, introducing appropriate semantic perturbation can significantly impact its quality predictions, while similar quality scores ensure that the misprediction is mainly influenced by the semantic perturbation rather than the quality perturbation.
@@
Figure 2: The overview of our QFM-IQM.We begin by extracting image semantic-related features from the student network. Simultaneously, we indirectly augment our training set by introducing a distillation framework and storing the image features and labels in pairs in temporary memory (Sec. 3.4). We use Semantic Noise Matching to select features in the temporary memory with similar quality scores but different semantic features as noise disturbance (Sec. 3.5) and add them to our original features as decoder input to predict image quality, forcing our model to learn to focus on quality-aware features (Sec. 3.6). During inference, scores are directly predicted using the encoder and decoder, while the above three modules are discarded to avoid any additional computational overhead.
@@ 2 Related Works
Image Quality Assessment (IQA) (Hu et al., 2021) is broadly classified into Full-Reference (FR) (Cao et al., 2022), Reduced-Reference (RR) (Tao et al., 2009), and No-Reference (NR) methods. NR-IQA assesses quality independently, offering significant application potential.
@@ 2.1 NR-IQA with Vision Transformer
Vision Transformer (ViT) (Dosovitskiy et al., 2021) showed promising results on several downstream vision tasks. There were mainly two types of ViT-based NR-IQA methods, including hybrid Transformer  (Golestaneh et al., 2022) and pure ViT-based Transformer  (Ke et al., 2021b; Yang et al., 2022; You & Korhonen, 2021). The hybrid architecture generally combined the CNNs with the Transformer, which was responsible for the local and long-range feature characterization, respectively. For instance, (Golestaneh et al., 2022) proposed to use the multi-scale features extracted from ResNet-50, which were fed to the transformer encoder to produce a non-local representation of the image. (Ke et al., 2021b) designed a multi-scale ViT-based IQA model to handle the arbitrary size of input images. Such a feature was initially designed to describe the image content, and thus the preserved features were mainly related to the higher-level visual abstractions which were not adequate in characterizing the quality-aware features(Qin et al., 2023).
@@ 2.2 NR-IQA with Contrastive Learning
A new learning paradigm, contrastive learning, has emerged for learning discriminative representations among samples (Caron et al., 2020; He et al., 2020). This method aims to learn an embedding space where similar samples are attracted, and dissimilar ones are repelled (Jaiswal et al., 2020; Zhao et al., 2023). Recently, contrastive learning in IQA has primarily focused on unsupervised approaches, as seen in key studies like QPT (Zhao et al., 2023), which introduced a self-supervised method with quality-aware contrastive learning for BIQA. Re-IQA (Saha et al., 2023) employed a mixture of experts to train separate encoders for content and quality features, and CONTRIQUE (Madhusudana et al., 2022) used a deep CNN with a contrastive pairwise objective for IQA. These methods have achieved promising results by leveraging self-supervised training, but the cost of pre-training models was often high. In contrast to contrastive learning aimed at constructing discriminative embedding spaces, we leverage adversarial learning (Goodfellow et al., 2014; Madry et al., 2017) to reduce the model’s sensitivity to semantic noise. This enhances the model’s ability to distinguish between images that are semantically similar but have different quality scores.
@@ 3.1 Preliminaries
In the context of No-Reference Image Quality Assessment (NR-IQA), we define some common notations. We use bold formatting to denote vectors (e.g., $\boldsymbol{x}$, $\boldsymbol{y}$) and matrices (e.g., $\boldsymbol{X}$, $\boldsymbol{Y}$). Specifically, $\boldsymbol{F}$ denotes the feature map of the network output, which includes the VIT’s class (CLS) token and patch embedding, and $\boldsymbol{y}$ denotes the quality score.
@@ 3.2 Overview
In this paper, we present Quality-Aware Feature Matching Image Quality Metric (QFM-IQM), a novel framework that can effectively distinguish between images with similar semantic features but different quality scores. As depicted in Fig. 2, QFM-IQM seamlessly integrates three main components: Semantic Noise Matching (SNM), Quality Consistency Constraint (QCC) and Distillation Label Extension (DLE). Initially, the input distorted image is fed into a transformer encoder to obtain feature $\boldsymbol{F}_{o}\in\mathbb{R}^{(M+1)\times D}$ from the final layer. Subsequently, the SNM module matches these features with those that have similar quality scores but different semantics (Sec. 3.5). Next, the matched semantic noise $\hat{\boldsymbol{F}}_{o}$ is introduced as feature perturbation into feature $\boldsymbol{F}_{o}$ to obtain feature $\mathcal{F}_{o}$, and the QCC module constraints between the quality prediction of the $\boldsymbol{F}_{o}$ and $\mathcal{F}_{o}$ to remain consistent, forcing the model to be robust to quality-unrelated semantic noise during adversarial learning (Sec. 3.6). Additionally, the DLE module uses knowledge distillation to generate pseudo-labels for unlabeled samples, thereby enriching the adversarial learning dataset of QCC (Sec. 3.4). These components collectively contribute to a more effective and accurate NR-IQA process. It’s worth noting that during the inference process, scores are predicted directly using the encoder and decoder, discarding the above three modules, and avoiding any additional computational overhead.
@@ 3.3 Architecture Design
Transformer Decoder. Previous works (Qin et al., 2023) found that CLS tokens cannot build an optimal representation for image quality. we utilize a quality-aware decoder that further interprets the CLS token. The decoder uses Multi-Head Cross-Attention (MHCA) with a set of queries $Q_{d}$, interacting with encoder outputs to calculate the final quality score $\boldsymbol{y}$. This method improves the learning and generalization capabilities of our NR-IQA model.
@@ 3.5 Semantic Noise Matching
Semantic Noise Feature Matching. This process involves selecting images with similar quality but different semantics to use as semantic noise features. Then, these noise features are employed as feature perturbations to disrupt semantics while largely preserving quality information. Finally, under the constraint of quality consistency, this effectively reduces sensitivity to semantic noise (referenced in Sec. 3.6). Fig. 3 overviews our Semantic Noise Matching (SNM) module. For a mini-batch of distorted input images, Temporary Memory $\mathcal{A}$ and $\mathcal{B}$ store collections of features and labels, then QFM-IQM leverages the SNM module to select noise samples that have similar quality scores but different semantic features from both labeled and unlabeled datasets. Specifically, the process begins by flattening the feature $\boldsymbol{F}_{o}$ of the input image, obtained from the encoder, into a feature vector $\boldsymbol{f}_{o}$. Subsequently, another feature vector $\hat{\boldsymbol{f}}_{o}$ is matched from the Temporary Memory $\mathcal{A}$ or $\mathcal{B}$ based on the matching strategy. The evaluation metric for the matching strategy is formulated as the product of feature similarity and score distance. Feature similarity is calculated using cosine distance, while score distance is determined using the $\mathcal{L}_{2}$ norm. The mathematical expression is as follows:
$\mathcal{S}((\boldsymbol{{f}}_{o},\rm{{y}}_{o}),(\hat{\boldsymbol{f}}_{o},\rm{\hat{{y}}_{o}}))=\frac{\boldsymbol{{f}}_{o}\cdot\hat{\boldsymbol{f}}_{o}}{\|\boldsymbol{{f}}_{o}\|\cdot\|\boldsymbol{\hat{{f}}_{o}}\|}\cdot\|\rm{{y}}_{o}-\rm{\hat{{y}}_{o}}\|.$ (1)
@@ 3.6 Quality Consistency Constraint
Existing studies have not addressed the challenge of differentiating between images with similar semantic features but different quality scores and eliminating quality-irrelevant features from high-level pre-trained features. To address this, we introduce the Quality Consistency Constraint (QCC) module to isolate quality-related attributes from an image’s semantic content. In contrast to previous contrastive learning approaches, this paper introduces a straightforward yet effective adversarial method to enhance a model’s robustness to features unrelated to image quality. This is accomplished by strengthening the consistency of the model’s predictions for image quality before and after introducing noise to distorted images. This method reduces the model’s sensitivity to semantic noise and compels it to exhibit robustness to subtle variations, ultimately increasing the separation between feature representations of noise samples. Concretely, we first extract the K features with the smallest evaluation metric calculated in the SNM module and incorporate them as feature perturbations into the original features. Since we perform feature selection twice in labeled dataset $\mathcal{A}$ and unlabelled dataset $\mathcal{B}$, the expression of this process is:
${\mathcal{F}_{o}}=(1-{\lambda_{1}}-{\lambda_{2}}){{{\boldsymbol{F}}}_{o}}+\frac{\lambda_{1}}{K}\sum_{j=1}^{K}{\boldsymbol{\hat{F}}^{\mathcal{A}}_{o,j}}+\frac{\lambda_{2}}{K}\sum_{j=1}^{K}{\boldsymbol{\hat{F}}^{\mathcal{B}}_{o,j}},$ (2)
where the weight $\lambda_{1}$ and $\lambda_{2}$ controls the strength of the perturbation. ${\boldsymbol{\hat{F}}^{\mathcal{A}}_{o,j}}$ and ${\boldsymbol{\hat{F}}^{\mathcal{B}}_{o,j}}$ represent the $j^{th}$ smallest matching feature from the datasets $\mathcal{A}$ and datasets $\mathcal{B}$. After that, we feed the feature $\mathcal{F}_{o}$ which is obtained by the Eq. 2 into the decoder to estimate the quality score $\boldsymbol{y}$. Notably, even after adding noise, the predicted quality is still supervised using the ground truth of the original image. By minimizing the prediction loss on samples with semantic noise, the model’s robustness to the semantic noise is improved.
@@ 4.4 Generalization Capability Validation
To assess the generalization capabilities of QFM-IQM, we conduct a series of cross-dataset validation experiments. In these tests, QFM-IQM trains on one dataset and then evaluates it on different datasets without adjustments to its parameters. The results, presented in Table 2, demonstrate QFM-IQM’s superior performance over SOTA models across six datasets. This includes notable gains on the LIVEC dataset and strong results on the KonIQ dataset. These results, benefiting from the noise-resistant training provided by the SNM and QCC, as well as the extensive semantic noise of various distortions provided by the DLE, underscore QFM-IQM’s exceptional ability to generalize.
@@ 4.5 Ablation Study
The core idea of our QFM-IQM is to improve the model’s ability to distinguish subtle quality differences by enhancing its robustness to semantic noise. Therefore, we explore the impact of Quality Consistency Constraint (QCC) operation on the model when compared with different datasets in Table 3. Let $\mathcal{A}^{QCC}$ and $\mathcal{B}^{QCC}$ denote the QCC operations applied only to the labeled dataset $\mathcal{A}$ and the unlabeled dataset $\mathcal{B}$, respectively. In the ablation experiments on the QCC module, we use the Semantic Noise Matching (SNM) module by default.
@@
Figure 5: Activation maps of DEIQT  (Qin et al., 2023) and QFM-IQM using Grad-CAM  (Selvaraju et al., 2017) on authentic dataset LIVEC. The scores below the first row of images represent Mean Opinion Scores (MOS). Our model pays more attention to image distortion regions (red boxes) and is less sensitive to semantic noise (yellow boxes), and correspondingly our image quality prediction ability is closer to the true value. Rows 1-3 represent input images, CAMs from DEIQT, and CAMs from QFM-IQM, respectively.
Effective on SNM and QCC. We extend our study to ablation experiments on SNM and QCC mechanisms, as shown in Table 4. Specifically, we undertake two sets of SNM ablation experiments: the first set, ${\text{SNM}}^{F}$, matches pairs of samples dissimilar in features, while the second set, ${\text{SNM}}^{Q}$, matches pairs similar in quality. Additionally, an extra ablation study without using the QCC involves matching noise samples $(\boldsymbol{x}_{1},\boldsymbol{y}_{1})$ and $(\boldsymbol{x}_{2},\boldsymbol{y}_{2})$ to a given pair $(\boldsymbol{x},\boldsymbol{y})$, where $\boldsymbol{y}$ is the quality label, adjusting the quality labels by a specific ratio ($\boldsymbol{\beta}_{1}=\boldsymbol{\beta}_{2}=0.1$) to incorporate noise sample labels, formulated as $\boldsymbol{y}=(1-\boldsymbol{\beta}_{1}-\boldsymbol{\beta}_{2}){\boldsymbol{y}}+{\boldsymbol{\beta}_{1}}{\boldsymbol{y}_{1}}+{\boldsymbol{\beta}_{2}}{\boldsymbol{y}_{2}}$.
@@
Table 5: Ablation experiments on the LIVEC and TID2013 datasets for unlabeled data selection in the DLE module.